Detemine the phase ahift of y = (3x-2pi)
Answer by lwsshak3(11628)

Detemine the phase ahift of y = (3x-2pi)
..
set y=0, then solve for x,
3x-2pi=0
3x=2pi
x=(2/3)pi
ans:
phase shift=(2/3)pi radians